# Uncertainty Quantification for Language Models: A Suite of Black-Box, White-Box, LLM Judge, and Ensemble Scorers

### v3

@@ 6 Conclusions
In this paper, we detail a framework for zero-resource hallucination detection comprised of various black-box UQ, white-box UQ, and LLM-as-a-Judge scorers. To ensure standardized outputs of the scorers, we transform and normalize scorers (if necessary) such that all outputs range from 0 to 1, with higher scores indicating greater confidence in an LLM response. These response-level confidence scores can be used for generation-time hallucination detection across a wide variety of LLM use cases. Additionally, we introduce a novel, ensemble-based approach that leverages an optimized weighted average of any combination of individual confidence scores. Importantly, the extensible nature of our ensemble means that practitioners can include additional scorers as new methods become available. To streamline implementation of the framework, all included scorers can be easily implemented using this paper’s companion Python toolkit, uqlm.
Our experimental evaluation of UQ-based scorers offers clear guidance for practitioners. Ensemble approaches consistently outperform individual methods for hallucination detection, with our findings strongly supporting use-case-specific customization rather than one-size-fits-all solutions. For those without token-probability access, NLI-based approaches typically provide the best black-box performance, though practitioners should be aware that all black-box methods struggle when response variation is limited. Importantly, gains from sampling additional responses diminish as the number of candidate responses increases, offering a practical deployment guideline that balances effectiveness with computational efficiency. Finally, our analysis revealed that a model’s accuracy on a specific dataset positively relates to its ability to judge responses to questions from that same dataset, providing a practical criterion for judge selection in evaluation frameworks.
@@ Appendix A LLM-as-a-Judge Prompt
To ensure a normalized confidence score consistent with the other scorers, we normalize the value returned by the LLM judge to be between 0 and 1. The capitalization and repeated instructions, inspired by Wang et al. (2024), are included to ensure the LLM correctly follows instructions.
@@ Appendix B Ensemble Tuning
We outline a method for tuning ensemble weights for improved hallucination detection accuracy. This approach allows for customizable component-importance that can be optimized for a specific use case. In practice, tuning the ensemble weights requires having a ‘graded’ set of $n$ original LLM responses which indicate whether a hallucination is present in each response. [^16] Grading responses may be accomplished computationally for certain tasks, e.g. multiple choice questions. However, in many cases, this will require a human grader to manually evaluate the set of responses. For a set of $n$ prompts, we denote the vector of original responses as $\mathbf{y}$
@@
$\hat{\mathbf{s}}(\mathbf{y};\tilde{\mathbf{Y}},\mathbf{w})=\begin{pmatrix}\hat{s}(y_{1};\tilde{\mathbf{y}}_{1},\mathbf{w})\\ \hat{s}(y_{2};\tilde{\mathbf{y}}_{2},\mathbf{w})\\ \vdots\\ \hat{s}(y_{n};\tilde{\mathbf{y}}_{n},\mathbf{w})\\ \end{pmatrix}$, (15)
$\hat{\mathbf{h}}(\mathbf{y};\tilde{\mathbf{Y}},\mathbf{w},\tau)=\begin{pmatrix}\hat{h}(y_{1};\tilde{\mathbf{y}}_{1},\mathbf{w},\tau)\\ \hat{h}(y_{2};\tilde{\mathbf{y}}_{2},\mathbf{w},\tau)\\ \vdots\\ \hat{h}(y_{n};\tilde{\mathbf{y}}_{n},\mathbf{w},\tau)\\ \end{pmatrix}$, (16)
@@
$\mathbf{h}(\mathbf{y})=\begin{pmatrix}h(y_{1})\\ h(y_{2})\\ \vdots\\ h(y_{n})\\ \end{pmatrix}$. (17)
@@
$\mathbf{w}^{*}=\operatorname*{arg\,max}_{\mathbf{w}\in\mathcal{W}}\mathcal{S}(\hat{\mathbf{s}}(\mathbf{y};\tilde{\mathbf{Y}},\mathbf{w}),\mathbf{h}(\mathbf{y})),$ (18)
@@
$\mathbf{\tau}^{*}=\operatorname*{arg\,max}_{\tau\in(0,1)}\mathcal{B}(\hat{\mathbf{h}}(\mathbf{y};\tilde{\mathbf{Y}},\mathbf{w}^{*},\tau),{\mathbf{h}}(\mathbf{y})),$ (20)
@@
$\mathbf{w}^{*},\tau^{*}=\operatorname*{arg\,max}_{\mathbf{w}\in\mathcal{W},\tau\in(0,1)}\mathcal{B}(\hat{\mathbf{h}}(\mathbf{y};\tilde{\mathbf{Y}},\mathbf{w},\tau),{\mathbf{h}}(\mathbf{y})),$ (21)
where $\mathcal{B},\hat{\mathbf{h}}$, $\mathbf{h},$ and $\mathcal{W}$ follow the same definitions as above.
Appendix C Additional Experiments: Number of Candidate Responses vs. Black-Box UQ Performance
@@
Overall, the results indicate that hallucination detection performance of the black-box scorers improves considerably with number of candidate responses $m$. For instance, hallucination detection AUROC of the various black-box scorers on GPT-4o responses on CSQA improve from 0.54-0.57 with $m=1$ to 0.75-0.8 with $m=15$. In the vast majority of our experimental scenarios, these performance improvements occur approximately monotonically with diminishing returns to higher $m$, consistent with findings from previous studies (Kuhn et al., 2023; Manakul et al., 2023; Lin et al., 2024; Farquhar et al., 2024). With the exception of BSC, this general trend is consistent in 21 of the 24 scenarios.
We observe two notable exceptions to this trend. First, BSC does not exhibit improved performance for higher $m$ in seven of the scenarios. This finding of overall less improvement in hallucination detection with higher $m$ for BSC is consistent with the experiments of Manakul et al. (2023) and is likely due to the token-wise nature of the comparisons.
Table 3: Average Exact Match Rate by LLM and Dataset
Model Used NQ-Open PopQA GSM8K SVAMP CSQA AI2-ARC
Gemini-1.5-Flash 0.81 0.79 0.83 0.96 0.99 1.00
Gemini-1.0-Pro 0.36 0.18 0.25 0.66 0.71 0.85
GPT-3.5 0.47 0.29 0.30 0.76 0.81 0.89
GPT-4o 0.44 0.35 0.63 0.91 0.90 0.81
Second, for Gemini-1.5-Flash responses on the AI2-ARC and CSQA datasets, we observe both poor black-box scorer performance (0.46-0.56) and a lack of improvement when increasing $m$. To investigate this finding, we computed the average exact match rate (EMR) across all 24 scenarios, with results presented in Table 3. Notably, Gemini-1.5-Flash exhibits remarkably high average EMR values on AI2-ARC (1.00) and CSQA (0.99), indicating almost no variation in the generated responses. This lack of diversity explains why increasing $m$ yields negligible improvement for these scenarios: additional samples provide virtually no new information.
More broadly, Gemini-1.5-Flash demonstrates consistently higher EMR across all benchmarks compared to other models. Accordingly, black-box hallucination detection is no better or worse on these Gemini-1.5-Flash responses compared to other LLM responses for the same dataset. This finding suggests that the effectiveness of black-box hallucination detection is fundamentally limited by response diversity and is consistent with findings by Kuhn et al. (2023), who find performance of various black-box scorers is considerably worse on smaller models that exhibit less response variation.

### v4

@@ 6 Conclusions
In this paper, we detail a framework for closed-book hallucination detection comprised of various black-box UQ, white-box UQ, and LLM-as-a-Judge scorers. To ensure standardized outputs of the scorers, we transform and normalize scorers (if necessary) such that all outputs range from 0 to 1, with higher scores indicating greater confidence in an LLM response. These response-level confidence scores can be used for generation-time hallucination detection across a wide variety of LLM use cases. Additionally, we introduce a novel, ensemble-based approach that leverages an optimized weighted average of any combination of individual confidence scores. Importantly, the extensible nature of our ensemble means that practitioners can include additional scorers as new methods become available.
Our experimental evaluation of UQ-based scorers offers clear guidance for practitioners. Ensemble approaches consistently outperform individual methods for hallucination detection, with our findings strongly supporting use-case-specific customization rather than one-size-fits-all solutions. For those without token-probability access, NLI-based approaches typically provide the best black-box performance. Importantly, gains from sampling additional responses diminish as the number of candidate responses increases, offering a practical deployment guideline that balances effectiveness with computational efficiency. Finally, our analysis revealed that a model’s accuracy on a specific dataset positively relates to its ability to judge responses to questions from that same dataset, providing a practical criterion for judge selection in evaluation frameworks.
The uqlm library implements all scorers presented and evaluated in this work. The repository is actively maintained. We welcome issues and pull requests and will continue to update integrations, add new scorers as research advances, and refresh examples as model and provider APIs change.
@@ Appendix A LLM-as-a-Judge Prompt
To ensure a normalized confidence score consistent with the other scorers, we normalize the value returned by the LLM judge to be between 0 and 1. The capitalization and repeated instructions, inspired by wang2024decodingtrustcomprehensiveassessmenttrustworthiness, are included to ensure the LLM correctly follows instructions.
@@ Appendix B Ensemble Tuning
We outline a method for tuning ensemble weights for improved hallucination detection accuracy. This approach allows for customizable component-importance that can be optimized for a specific use case. In practice, tuning the ensemble weights requires having a ‘graded’ set of $n$ original LLM responses which indicate whether a hallucination is present in each response. [^19] Grading responses may be accomplished computationally for certain tasks, e.g. multiple choice questions. However, in many cases, this will require a human grader to manually evaluate the set of responses. Given a set of $n$ prompts, denoted as $\mathbf{x}$:
$\mathbf{x}=\begin{pmatrix}x_{1}\\ x_{2}\\ \vdots\\ x_{n}\\ \end{pmatrix}$, (13)
denote corresponding correct reference answers as $\mathbf{y}^{*}$,
$\mathbf{y}^{*}=\begin{pmatrix}y_{1}^{*}\\ y_{2}^{*}\\ \vdots\\ y_{n}^{*}\\ \end{pmatrix}$, (14)
original LLM responses as $\mathbf{y}$,
@@
$\hat{\mathbf{s}}(\mathbf{y};\tilde{\mathbf{Y}},\mathbf{x},\mathbf{w})=\begin{pmatrix}\hat{s}(y_{1};\tilde{\mathbf{y}}_{1},x_{1},\mathbf{w})\\ \hat{s}(y_{2};\tilde{\mathbf{y}}_{2},x_{2},\mathbf{w})\\ \vdots\\ \hat{s}(y_{n};\tilde{\mathbf{y}}_{n},x_{n},\mathbf{w})\\ \end{pmatrix}$, (17)
$\hat{\mathbf{h}}(\mathbf{y};\tilde{\mathbf{Y}},\mathbf{x},\mathbf{w},\tau)=\begin{pmatrix}\hat{h}(y_{1};\tilde{\mathbf{y}}_{1},x_{1},\mathbf{w},\tau)\\ \hat{h}(y_{2};\tilde{\mathbf{y}}_{2},x_{2},\mathbf{w},\tau)\\ \vdots\\ \hat{h}(y_{n};\tilde{\mathbf{y}}_{n},x_{n},\mathbf{w},\tau)\\ \end{pmatrix}$, (18)
@@
$\mathbf{h}(\mathbf{y};\mathbf{y}^{*},\mathbf{x})=\begin{pmatrix}h(y_{1};y_{1}^{*},x_{1})\\ h(y_{2};y_{2}^{*},x_{2})\\ \vdots\\ h(y_{n};y_{n}^{*},x_{n})\\ \end{pmatrix}$, (19)
@@
$\mathbf{w}^{*}=\operatorname*{arg\,max}_{\mathbf{w}\in\mathcal{W}}\mathcal{S}(\hat{\mathbf{s}}(\mathbf{y};\tilde{\mathbf{Y}},\mathbf{x},\mathbf{w}),\mathbf{h}(\mathbf{y};\mathbf{y}^{*},\mathbf{x})),$ (20)
@@
$\mathbf{\tau}^{*}=\operatorname*{arg\,max}_{\tau\in(0,1)}\mathcal{B}(\hat{\mathbf{h}}(\mathbf{y};\tilde{\mathbf{Y}},\mathbf{x},\mathbf{w}^{*},\tau),{\mathbf{h}(\mathbf{y};\mathbf{y}^{*},\mathbf{x}))},$ (22)
@@
$\mathbf{w}^{*},\tau^{*}=\operatorname*{arg\,max}_{\mathbf{w}\in\mathcal{W},\tau\in(0,1)}\mathcal{B}(\hat{\mathbf{h}}(\mathbf{y};\tilde{\mathbf{Y}},\mathbf{x},\mathbf{w},\tau),{\mathbf{h}(\mathbf{y};\mathbf{y}^{*},\mathbf{x}))},$ (23)
where terms follow the same definitions as above.
Appendix C Response Grading
For each task-type, the LLM is instructed to output its response in a specific form. We therefore instantiate $h(\cdot)$ with task-specific graders aligned to that form. Unless noted, grading is case-insensitive and trims whitespace. [^20] These automatic graders may not be perfectly accurate in every instance; residual errors can arise when model outputs deviate from the specified response format. Their effectiveness therefore depends on instruction adherence, and they should be viewed as practical proxies for answer-key–based factual correctness.
Math (numeric answer only).
For math questions, the LLM is provided the following instruction: “Return only the numerical answer with no additional text.” Let $y_{i}$ be the model output. Let $\text{int}(y_{i})$ extract the leading integer substring from $y_{i}$ if present; otherwise return $\varnothing$. Let $y_{i}^{*}\in\mathbb{Z}$ be the correct integer answer.
$h_{\text{math}}(y_{i};y_{i}^{*})=\begin{cases}0&\text{if }\text{int}(y_{i})=y_{i}^{*},\\ 1&\text{otherwise.}\end{cases}$
Non-numeric or missing integer outputs are graded as incorrect.
Multiple choice (letter only).
For multiple choice questions, the LLM is provided the following instruction: “Return only the letter of the response with no additional text or explanation.” Let $\mathrm{norm}_{mc}(y_{i})\in\{A,B,C,D,E\}$ denote response $y_{i}$ after normalization to uppercase and trimming. Let $y_{i}^{*}\in\{A,B,C,D,E\}$ be the correct letter.
$h_{\text{mc}}(y_{i};y_{i}^{*})=\begin{cases}0&\text{if }\mathrm{norm}_{mc}(y_{i})=y_{i}^{*},\\ 1&\text{otherwise.}\end{cases}$
Any invalid response based on the provided instruction, i.e. $\mathrm{norm}_{mc}(y_{i})\notin\{A,B,C,D,E\}$, is graded as incorrect.
Short-answer (keyword match).
Lastly, for short-answer questions, the LLM is instructed as follows: “Return only the answer as concisely as possible without providing an explanation.” Let $\mathbf{y}_{i}^{*}=\{y_{i1}^{*},\dots,y_{iA}^{*}\}$ be the set of acceptable answers. [^21] The short-answer datasets contain multiple versions of the same correct answer to allow for phrasing flexibility. For example, acceptable answers to the question “What is Tobias Lindholm’s occupation?” include {“film director”, “movie director”, “director”, “motion picture director”, “screenwriter”, “scenarist”, “writer”, “screen writer”, “script writer”, “scriptwriter”}. Define a normalization function $\mathrm{norm}_{sa}(y_{i})$ that lowercases and trims $y_{i}$. Let $\mathrm{contains}(u,v)$ indicate that string $u$ contains string $v$ as a contiguous substring.
$h_{\text{sa}}(y_{i};\mathbf{y}_{i}^{*})=\begin{cases}0&\text{if }\exists\text{ }y_{i}^{*}\in\mathbf{y}_{i}^{*}\ \text{s.t. }\mathrm{contains}(\mathrm{norm}_{sa}(y_{i}),\mathrm{norm}_{sa}(y_{i}^{*})),\\ 1&\text{otherwise.}\end{cases}$
If output includes additional text, grading is based on the presence of any acceptable answer string.
Appendix D Additional Experiments: Number of Candidate Responses vs. Black-Box UQ Performance
@@ Appendix C Response Grading
Overall, the results indicate that hallucination detection performance of the black-box scorers improves considerably with number of candidate responses $m$. For instance, hallucination detection AUROC of the various black-box scorers on GPT-4o responses on CSQA improve from 0.54-0.57 with $m=1$ to 0.75-0.8 with $m=15$. In particular, these performance improvements occur approximately monotonically with diminishing returns to higher $m$, consistent with findings from previous studies (kuhn2023semanticuncertaintylinguisticinvariances; manakul2023selfcheckgptzeroresourceblackboxhallucination; lin2024generatingconfidenceuncertaintyquantification; Farquhar2024). This general trend is consistent across all 24 scenarios, with only a few exceptions, notably BSC for GPT-4o responses on GSM8k, and BSC, NCS, and EMR for GPT-4o responses on AI2-ARC.